1. Sam can brew 5 gallons of beer in an hour or he can make 4 pizzas in an hour. Ben can brew 7 gallons of root beer in an hour or he can make 5 pizzas in an hour. Who has an absolute advantage in making pizza sam/ben? who has a comparative advantage in making pizza? sam/ben

2.Latasha can read 30 pgs of economics in an hour. She can also read 20pgs of sociology in an hour. She spends 4 hours per day studying. Her opportunity cost of reading 10pgs of sociology is ________pgs of economics. Her opportunity cost of reading 18 pgs of economics is _____pgs of sociology. enter whole numbers

3.two countries produce two goods and they each have 3 millions hour of labor. consider the table that shows the quantity produced per labor hour.

shirt corn
alpha 4 12
beta 2 10

if country specializes in the good for which it has a comparative advantage then in total the two countries can together produce ____ million shirts and ______ corn. enter whole numbers
Answer #1

1. Sam can brew 5 gallons of beer in an hour or he can make 4 pizzas in an hour. Ben can brew 7 gallons of root beer in an hour or he can make 5 pizzas in an hour.

Ben has an absolute advantage in making pizza (higher productivity) and Sam has a comparative advantage in making pizza (lower opportunity cost for 1 pizza (Sam) = 5/4 = 1.25 gallons of beers)

2. Latasha can read 30 pgs of economics in an hour. She can also read 20 pgs of sociology in an hour. She spends 4 hours per day studying.

Her opportunity cost of reading 10 pgs of sociology is = 30/2 = 15 pgs of economics. Her opportunity cost of reading 18 pgs of economics is = 20/30 x 18 = 12 pgs of sociology.

3.

After specialization:

Shirts = 4 x 3 = 12 mn shirts (alpha would specialize in shirt production, lower opportunity cost = 12/4 = 3 units of corn)

Corn = 10 x 3 = 30 mn